Personal tools
You are here: Home Virgo Data Analysis VDAS vdas_20120503
Log in

Forgot your username or password?
Document Actions


by Marie Anne Bizouard last modified 2012-05-03 18:07


  1. Announcements
  2. VSR4 RDS transfer [Livio]
  3. BuffervXX cleanup at Cascina [Didier] [status page]  [allocations]
  4. ER2 preparation:
    • LSC (LDG) software to be installed at Cascina [Frederique]
    • What else?
    • Update on the current situation and special requests


Attendance: Elena, Alberto, Didier, Florent, Franco, Frederique, Nicolas, Benoit


3. BuffervXX:

Didier explains the situation. Monitoring page. Old data could be deleted?

MAB: procdata buffers are not backuped. Could be a problem because we have databases, triggers and important data dored in this area. Need to discuss this with EGO computing. May need to change this strategy. About size increase, MAB feeling is that maybe a factor 2-3 will be needed in 3-4 years from now. Need to know if the current architecture can be scaled.

4 ER2 preparation:

Demand for LDG software installation. Use of private version. MBTA send triggers to GRACEdb. Made test. Pb with security. Need to disable security at UWM. Can't continue like that. Need to be compliant with security rules. Need to install software as root. Need of SL6?

--> Franco will investigate the problem. Install LDG and report to VDAS/Frederique

Data generation:

  • Virgo: VSR2 real data rescaled.
  • LIGO: Gaussian noise + auxiliary channels + coupling simulated. 

Elena:  will there be also non linear coupling?

Nicolas: such thing will be discussed at the DASWG F2F meeting.

-->need to generate DQs to be sent to segdb. Could use the xml files. Need to change the GPS shift.

Question: do we want to feed VDB?

Question: do we know if the recoloring will affect the DQ flagging?

Nicolas: we should make a test with 1 week of data.

Frederique: h(t) is going to be generating from old h(t).

Glitchiness level of LIGO data? DA are asking lots of glitches.


Online processing:

  • Florent: Omicron on h(t).
  • Elena: non linear coupling. Need to coordinate with Gianluca and Marco. D-NMAPI @ Cascina
  • Alberto: NoEMi run on LIGO data. Need to check availability with LIGO


Low latency data transfer:

Test are running right now. Gaussian noise h(t) + adding injections. 1 month of test.


1 channel for all blind injections.

1 different channel for each unblind type of injections (burst, cbc)

Envelope opening: 2 weeks after the end.

Need to know when the injection is on? Flag added to h(t). Need to check.

All injections generated in advance.


Who is doing what:

Benoit : data generation + injections + low latency transfer

Didier: xml files generation + transfer to segdb

Frederique, Florent, Elena, Alberto: online processing



Franco: SL6 basic software versions + latest version of CMT.

pb with SL6/NSF. Showstopper for SL6